Your Algebra 1A teacher is 21 years older than you. The sum of your teacher and your age is 49.
Answer by Paul(988)

21+x+x=49
2x+21-21=49-21
2x=28
x=14
14+21=35
hence, you are 14 and your teacher is 35, years old.